OpenStudy (anonymous):

–10 + 3x + 6x = –10 A. –9 B. –10 C. 0 D. 9

OpenStudy (anonymous):

-10 + 3x + 6x = -10 -10 + 10 + 9x = -10 + 10 9x = 0 x=0

OpenStudy (anonymous):

thx fried